East Dome Status   2021-May-24 09:13:17 UTC



Device
Mode
Date/time(UTC) last changed Position
   Telescope Parked 2021-Apr-30    15:48:07
   Dome Shutter Closed (DDW) 2021-Apr-30    15:58:36
   Dome Shutter Not-Closed (optical sensor) 2021-Feb-08    06:54:01
   Dome Mode Not tracking 2021-Apr-25    21:06:02
Power Status
Telescope
OFF
2021-Apr-30    15:50:01
Dome
ON
2021-Apr-03    15:09:25
CCD
OFF
2021-May-02    11:25:25
Heating Strip
ON
2021-Mar-17    17:44:39
Dome LED Lights
OFF
2021-Apr-25    21:10:32